• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name M Friend

    Log entry time 08:00:48 on October12,2009

    Entry number 294529

    keyword=Owl Shift Summary

    00:00 Beam down due to MCC ioc problems.

    00:12 Requested an access for Compton FADC trigger threshold change. Perfect since they were going down there anyway to do a hard reboot of the ioc.
    00:36 MCC has finished the ioc reboot

    00:47 Tune beam
    00:50 5uA beam for spot++; the left arm is having DAQ problems and requires rebooting the ROCs (right arm can't be used for spot++ because of a bpm problem). Spot++ looks fine.
    00:57 Up to 100uA. Beam very trippy.

    01:25 MCC made a change which fixed the Compton rates, but they had to back out because it caused the beam to trip too much.

    Compton tuning by MCC is ongoing.

    5:28 MCC giving up on compton tuning.

    6:51 Change IHWP to IN, begin slug 22


    ~Production run 13183, 6.4k good pairs
    ~Pedestal run 13184
    Production run 13185, 20.9 good pairs
    Production run 13187, 41.2k good pairs
    Production run 13188, 25.4k good pairs
    Production run 13189, 41.4k good pairs
    Production run 13191, 40.6k good pairs
    Production run 13192, 40.9k good pairs
    IHWP is IN
    Production run 13194, 43.4 good pairs
    Production run 13195, ongoing